OpenStudy (anonymous):

Which of the following equations have the same solution? Solve all three equations separately and explain each step.

OpenStudy (anonymous):

One moment, putting it up

OpenStudy (anonymous):

I. 2x + 4 = 3x – 11 II. 3x + 3 = 51 III. \[\frac{ 2x + 1 }{ 3 } = \frac{ x + 6 }{ 2 }\]

OpenStudy (anonymous):

I. Solve for x: 2 x+4 = 3 x-11 Subtract 3 x from both sides: (2 x-3 x)+4 = (3 x-3 x)-11 2 x-3 x = -x: -x+4 = (3 x-3 x)-11 3 x-3 x = 0: 4-x = -11 Subtract 4 from both sides: (4-4)-x = -11-4 4-4 = 0: -x = -11-4 -11-4 = -15: -x = -15 Multiply both sides of -x = -15 by -1: (-x)/(-1) = 15 (-1)/(-1) = 1: Answer x = 15

OpenStudy (anonymous):

II. Solve for x: 3 x+3 = 51 Subtract 3 from both sides: 3 x+(3-3) = 51-3 3-3 = 0: 3 x = 51-3 51-3 = 48: 3 x = 48 Divide both sides of 3 x = 48 by 3: (3 x)/3 = 48/3 3/3 = 1: x = 48/3 The gcd of 48 and 3 is 3, so 48/3 = (3\[Times]16)/(3\[Times]1) = 3/3\ \[Times]16 = 16: Answer: x = 16

OpenStudy (anonymous):

III. Solve for x: (2 x+1)/3 = (x+6)/2 Multiply both sides by 6: (6 (2 x+1))/3 = (6 (x+6))/2 6/3 = (3*2)/3 = 2: 2 (2 x+1) = (6 (x+6))/2 6/2 = (2*3)/2 = 3: 2 (2 x+1) = 3 (x+6) Expand out terms of the left hand side: 4 x+2 = 3 (x+6) Expand out terms of the right hand side: 4 x+2 = 3 x+18 Subtract 3 x from both sides: (4 x-3 x)+2 = (3 x-3 x)+18 4 x-3 x = x: x+2 = (3 x-3 x)+18 3 x-3 x = 0: x+2 = 18 Subtract 2 from both sides: x+(2-2) = 18-2 2-2 = 0: x = 18-2 18-2 = 16: Answer: x = 16

OpenStudy (anonymous):

II and III have the same solution.

OpenStudy (anonymous):

@half I see... thank you for the help!